A cricketer can throw a ball to maximum horizontal distance of 150m .How high above the ground can the cricketer throw the same ball
1
Expert's answer
2019-01-24T10:25:32-0500

To reach the maximum horizontal distance a racketeer should throw a ball with angle

"45^\circ"

to horizontal. So

"L=\frac{v_0^2\cdot sin2\alpha}{g}=\frac{v_0^2}{g}"

To reach maximum heigh the angle must be

"90^\circ"

:

"H=\frac{v_0^2\cdot sin^2\alpha}{2g}=\frac{L}{2}=75 m"